Questions 5 – 7 use the following setup.

Summit Systems will pay a dividend of $1.50 next year. If you expect Summit’s dividend to grow by 6% per year and its required return is 11%, what is its price per share today?

Under which of the following scenarios will you be willing to pay a higher price than what you’ve solved for in Question 5?

Group of answer choices

A. The dividend will be $1.20 next year.

B. The required return is greater than 11%.

C. The growth rate is 8% per year.

D. Both (B) and (C) are correct.

What will its stock price be in four years?

Group of answer choices

A. $35.73

B. $37.87

C. $31.80

D. $45.54
Answer #1

5. Price per share today =Dividend next year/(Required Rate-Growth) =1.50/(11%-6%) =30

6. Option c is correct option The growth rate is 8% per year.
Other options are incorrect because lower the dividend payout and higher the required rate , lower is the price of bond.

7. Stock price after 4 years =Current price*(1+growth)^4 =30*(1+6%)^4 =37.87
(Option b is correct option)